Thirty-four automobiles were tested for fuel efficiency (in miles per gallon). The following frequency distribution was obtained. Find the variance and standard deviation. Round your answers to one decimal place. Class boundaries 7.5-12.5 12.5-17.5 17.5-22.5 22.5-27.5 27.5-32.5 Frequency 14 Download data Variance- Standard deviation =
Answer #1
Class interval Class mark Frequency
7.5-12.5 10 3
12.5-17.5 15 9
17.5-22.5 20 14
22.5-27.5 25 7
27.5-32.5 30 1

Mean = Sum of (Class mark x frequency) / Sum of frequency

= (10x3 + 15x9 + 20x14 + 25x7 + 30x1) / (3+9+14+7+1)

= 650/34

= 19.1

Variance = Sum of [(Class mark - mean)2 x frequency] / Sum of frequency

= [(10 - 19.1)2x3 + (15-19.1)2x9 + (20-19.1)2x14 + (25-19.1)2x7 + (30-19.1)2x1]/34

= 773.5/34

= 22.8

Standard deviation = \sqrt{22.8}

= 4.8